# Space-time Peer-to-Peer Distribution of Multi-party Entanglement for Any Quantum Network

### v1

@@ Abstract.
Graph states are a class of important multiparty entangled states, of which bell pairs are the special case. Realizing a robust and fast distribution of arbitrary graph states in the downstream layer of the quantum network can be essential for further large-scale quantum networks. We propose a novel quantum network protocol called P2PGSD inspired by the classical Peer-to-Peer (P2P) network to efficiently implement the general graph state distribution in the network layer, which demonstrates advantages in resource efficiency and scalability over existing methods for sparse graph states. An explicit mathematical model for a general graph state distribution problem has also been constructed, above which the intractability for a wide class of resource minimization problems is proved and the optimality of the existing algorithms is discussed. In addition, we proposed the spacetime quantum network inspired by the symmetry from relativity for memory management in network problems and used it to improve our proposed algorithm. The advantages of our protocols are confirmed by numerical simulations showing an improvement of up to 50% for general sparse graph states, paving the way for a resource-efficient multiparty entanglement distribution across any network topology.
@@ 2.2. Principles of P2P Graph State Distribution
Here, $G_{N}=(V_{N},E_{N})$ is an undirected graph representing the topology of the network with $V_{N}$ denoting the set of nodes and $E_{N}$ denoting the set of channels. $W_{c}:E_{N}\rightarrow\mathbbm{N}^{+}$ is a map that represents the width of the channel such that each channel $c\in E_{N}$ can establish $W_{c}(c)$ bell pairs per cycle time. And $W_{m}:V_{N}\rightarrow\mathbbm{N}$ is a map that represents the long-term memory available for each node.
@@ 2.3. Problem Formulation
An One-shot Bell Pair Strategy $b_{1}$ is a map $s_{1}:\vec{E}_{N}\times V_{S}\rightarrow\{+1,0,-1\}$ with 00 representing that the channel is not used by the corresponding vertex and it fulfills the following flow conditions:
@@
    $\exists n_{j}(\neq n_{i})\in N_{v_{s}}^{\mathscr{b}_{N}},\,b_{k}\in\mathscr{b}_{N},\vec{c}_{ij}=(c_{ij},n_{j},n_{i})\in\vec{E}_{N},\mathrm{s.t.}\,b_{i}(\vec{c}_{ij},v_{s})>0$.
@@
A $N$-Shot Bell Pair Flow Strategy $[\mathscr{s}_{N}]$ is an equivalent class defined by the following equivalent relation for $N$-shot Trial Solution as
$\mathscr{b}_{N}\sim\mathscr{b}^{\prime}_{N}\Leftrightarrow\sum_{b_{i}\in\mathscr{s}_{N}}{b_{i}|(\vec{e},v_{s})|}=\sum_{b^{\prime}_{i}\in\mathscr{s}^{\prime}_{N}}{|b^{\prime}_{i}(\vec{e},v_{s})|},\,\forall v_{s}\in V_{S},\,\forall\vec{e}\in\vec{E}_{N}.$
@@ 2.4. Optimizing Objectives
The first metric is simply the number of cycles required, and it is also the time consumption to execute such a solution. The second is frequently considered in the previous literature (Cuquet and Calsamiglia, 2012; Pirker et al., 2018; Meignant et al., 2019; Fischer and Towsley, 2021; Xie et al., 2021; Fan et al., 2024), while the Long-term Memory Consumption, as we explained in the previous subsection, is just the cumulative memory (Alwen and Serbinenko, 2015; Beame and Kornerup, 2023) and can be considered as a counterpart of the second. We remark that some qubits during execution can be measured instantaneously, while some other qubits will be required to be kept to the next shot. The latter must be kept for at least one cycle time, where the consumption of quantum memory is necessary and the usage is being counted by our defined metric.
@@ 2.5. Complexity Analysis
The aforementioned theorem is important in the sense that it forces us and also rationalizes it to pursue a heuristic algorithm that consumes low resources with relatively low runtime, since it is nearly impossible to have a scalable and practical one that converges to the global minimum unless $P=NP$. However, we get a tight upper bound for #$\#\mathscr{S}$ in the case where there is unlimited memory. Here, ”tight” means that we construct an instance $\mathscr{P}$ that requires at least the upper bound number of shots to distribute using any algorithm. This result will help us analyze the performance of the proposed algorithms in the worst case and also ensures that the algorithms obtained are within a polynomial approximation (linear to optimize #$\#\mathscr{S}$) and thus reliable.
@@ 3.1. Network Model and Routing Protocol
    The high quality bell pair is created in a heralded manner (it can be an abstraction of the physical devices that generate high quality bell pairs from multiple tries and entanglement purification, including the quantum repeater(Bennett et al., 1996; Bernien et al., 2013; Pirandola et al., 2017)). A channel of width $W_{c}(e)$ with $e\in E_{N}$ is capable of creating at most $W_{c}(e)$ bell pairs within a time interval, each with a probability of $P_{c}(e)$.
@@
    All single-qubit operations are assumed to be deterministic, but the only two-qubit operation, $CZ$ operation, is considered probabilistic with probability $P_{s}$ but in a heralded way (Li et al., 2020; Sung et al., 2021, 2021).
@@ 3.2. Modified Graph State Transfer Algorithm
Figure 3. Illustration for MGST algorithm. (a) The constructed flow graph with the subgraphs inside the red shadow box represent one of the copies of $G_{N}$ labelled as $G_{N}^{i},\,i=1...k$ together with virtual nodes (larger circles) and virtual arcs (dotted arrows) with the numbers in different colors representing the width of corresponding channel. (b) An example of $3$-shot solution of MGST. The nodes with multiple colors is the root. The paths in different colors represent the edge-disjoint spacetime path from $\{(\alpha_{D}(v_{s}),4\}$ to the root node $(n_{\mathrm{root}},1)$. The blue dotted arcs show the possible recovery possible if the current path fails.
@@
The GST algorithm proposed by Fischer et al. (Fischer and Towsley, 2021) is an efficient algorithm but it only gives us an equivalent class $[\mathscr{b}_{N}]$ that solves $\mathscr{P}$. We need to pick an element $\mathscr{b}_{N}$ from $[\mathscr{b}_{N}]$ as well as find a $\mathscr{m}_{n}$ to form the solution $\mathscr{S}_{N}=\{\mathscr{b}_{N},\mathscr{m}_{N}\}$ so that our network can execute the solution. One can certainly use the flow decomposition theorem to decompose $[\mathscr{b}_{N}]$ into a set of paths; however, it is unknown whether such paths can be executed in $N$ shots without conflicting or taking apart specific paths into several shots (which we call it simply executable) is unknown. In the worst case, we may have to consume a lot of memory to execute a single path parts-by-parts across multiple shots, which is definitely unwanted. Hence, how to decompose the flow into a $N$-shot simple executable path set is not a trivial problem. We propose the following novel construction of the flow graph to find the $|V_{N}|$ edge-disjoint paths from a given root $n_{\mathrm{root}}\in V_{N}$ executable in $k$ shots:
@@
However, there is one question left to be solved: Will finding such a simple executable path set result in the use of more shots than the flow solution given by GST. This question is answered by the following theorem proven in Appendix. A.1.
@@ 3.3. Peer-to-Peer Graph State Distribution Algorithm
Figure 4. Illustration for P2PGSD algorithm. (a) When the path between $n_{v_{s_{1}}}$ (yellow) and $n_{v_{s_{2}}}$ (blue) is first found, the assignment of the bell pairs (yellow & blue) is not determined. After the new path to $n_{v_{s_{3}}}$ (green node) is found from an ’unfixed’ node along Path 1, part of the Path is fixed and it is split into the new Path 1 and Path 2, while Path 3 is not fixed (yellow & green). (b) The P2PGSD solution under Minimum memory strategy. The half-and-half colors indicate the node that implement the edge for the corresponding vertices. (c) The P2PGSD solution under Standard memory strategy. (d) The P2PGSD solution under Maximum memory strategy. This might be helpful for later trials (dotted arcs) in case the previous trial fails.
@@
We omitted the proof of the correctness for this simple modification version of Dijkstra, while we remark that the correctness of this algorithm relies on the assumption that in the current network, no path from other $n_{v_{s_{1}},i}/n_{v_{s_{2}},j}$ to other $n_{v_{s_{1}},i^{\prime}}/n_{v_{s_{2}},j^{\prime}}$ with smaller cost exits is available. The modified version of the Dijkstra algorithm has a time complexity of $\mathscr{O}(|E_{N}||V_{N}|)$ where the additional $|V_{N}|$ factor comes from the fact that we need to calculate the metric for the whole path at each time to allow a more general metric beyond the addable one, though we can turn our metric into the latter in this work.
@@
    Maximum: All the nodes’ ancillary vertices will be kept to the next shot.
@@ 3.4. Spacetime Peer-to-Peer Graph State Distribution Algorithm
Figure 5. Illustration for STP2PGSD algorithm. (a) A no-plan-ahead solution compared to (b) a solution without such limitation, where the parallel arcs in the shot $k$ of the lower plane indicate the violation of such limitation. (b) STP2PGSD directly finds a path in the spacetime network and overcomes the limitation of no-plan-ahead.
@@
The pseudocode for the STP2PGSD algorithm is shown in Alg. 3 and the runtime of STP2PGSD is upper bound by $o(|V_{S}|^{4}|E_{S}||V_{N}||E_{N}|\log|V_{S}|)$.
@@ 3.5. Adaptive Protocol
At each shot, only the plan for the first shot given by the algorithm will be collected and executed by the network. After that, the execution results for the current shot will be collected in the synchronization phase and the distribution task $\mathscr{P}$ will be modified accordingly. Then the modified distribution task together with the available resources from the last shot will be submitted to the algorithm in the next cycle to perform an adaptive distribution strategy. This mechanism offers strong flexibility for the graph state distribution in a probabilistic network and also inherits a sense of recovery mechanism using the link resources from the previous shot, which will be beneficial to scaling issues when the probabilities of success for the links in the network are low. However, such an adaptive protocol will require some modifications to the original algorithm to take care of the resources from the previous shot, which is easy and trivial for the first two algorithms but requires careful design for the ST-P2PGSD algorithm. The further details of the adaptive protocol will be left for discussion in the Appendix. C.
@@ 4.1. Simulation Setup
The proposed network models as well as the algorithms are implemented on a custom-built simulator in Python with topology generation and common subroutines adapted from the NetworkX package (Hagberg et al., 2008). As we have mentioned in Sec. 3.1, the protocols presented can be considered a network protocol that regards general graph state as a basic connection instead of bell pairs. Thus, our simulations are not built upon other proposed protocols, but rather be implemented from the relatively basic level. The source codes as well as the data for the numerical evaluation can be found in the corresponding open repository (Yuexun Huang, 2024).
@@
Multiple types of graph state are being considered including the Erdos-Renyi graph (ERDdS and R&wi, 1959), star graphs, Prufer tree graphs (Wang et al., 2009), grid graphs, etc., with some of them being generated randomly when possible. After the graph state topology is generated, the vertices are randomly assigned to the network nodes to implement the assignment map $\alpha_{D}$.
@@ 4.2. Metrology
The simulation is designed to evaluate the time shots, #$\#\mathscr{S}$, and cumulative memory consumption, $\#\mathscr{M}$, defined in Sec. 2.2 as well as the computation runtime for distributing various types of graph states (random or fixed) on a random network under the algorithms proposed in this work as well as the one modified from previous work (MGST). We do not focus on bell pair usage, $\#\mathscr{B}$, since it is not our main optimization goal and it can be shown that it is not compatible with optimization of the previous two resources and requires a completely different design strategy for the algorithms.
@@ 4.3. Results Analysis
The general performance of protocols based on 3 main algorithms together with two additional variants for P2PGSD and ST-P2PGSD are shown in Fig. 7, where we conclude that the developed P2P series algorithms demonstrate a huge advantage over the center-based MGST algorithm in distributing the widely used tree graphs and grid graphs in the sense of shot usages and cumulative memory usage. The resource consumption difference grows when the number of vertices of the graph state increases, which supports our statement that many resources can be wasted if the graph state topology is not considered for distribution. Furthermore, P2PGSD with maximum memory strategy consumes more memory than P2PGSD while producing almost the lower shot consumption, which illustrates the idea of trading memory for faster distribution. However, one also notices that the ST-P2PGSD consumes memory comparable to P2PGSD, but also a shot consumption that is sometimes lower than P2PGSD-Max. This is due to the more efficient memory planning enabled by the ST network. We also investigate the performance of the algorithms under extremely limited memory as shown in Fig. 7(c) where each node only holds the amount of long-term memory that is equal to its assigned vertices to fulfill the minimum requirement to distribute such a graph state (while one of the nodes in the network is chosen to possess $2|V_{S}|$ memory in order for MGST to be executed). In this case, MGST consumes significantly more resources than the P2PGSD algorithms series and also has huge variance due to the available center node choice. In addition, the performance of various P2PGSD algorithms is drawn to a common line since there is not much room for memory optimization.
@@
Figure 7. Shots (bar) and cumulative memory (scatter) usage of protocols based on various algorithms versus the number of vertices in certain types of graph state. P2PGSD adapts the standard memory strategy while P2PGSD-Max adapts the maximum memory strategy. The shaped area and the error bar represent the standard deviation of the data. The metric of standard ST-P2PGSD is given by Eq. (11) while the ”Factor” version is given by Eq. (12). (a) Random Prufer tree graphs. (b) Grid graphs. (c) Grid graphs while the performance is evaluated under extremely limited memory. (d) Star graphs. (e) Bell pairs (regular 1) graphs with cell topology network (illustrated in subfigure). (f) Erdos-Renyi with average edge probability $0.1$.
@@
Advantageous tasks. Three different main algorithms establish significant advantages with respect to distribution tasks with certain topology characteristics. Fig. 7 (d)-(f) shows the typical performance comparison between the algorithms on the presentative tasks. As our initial inspiration for this work, the P2PGSD series algorithms demonstrate extremely low resource consumption when the state being distributed is a star graph (GHZ state), as shown in Fig. 7 (d). One observes that shot usage as well as cumulative memory consumption hardly grows under the P2PGSD series algorithms, in contrast to MGST. This is a straightforward consequence since distributing such a state only requires the assigned node to be connected in tree form by the successful bell pair and the resource consumption saturates if the network size is unchanged due to the deterministic local operation. The spacetime version algorithms do not show any advantages in the previous case since these tasks admit a one-shot solution and do not need memory optimization, but one observes from Fig. 7 (b) that they consume remarkably fewer resources when the network suffers from serious discontinuous bottleneck channels as the case of the so-called ”cell topology network” shown in Fig. 7 (g). The ST-network method makes it rather easy and convenient to deal with this problem and generate an efficient memory management plan. Finally, the centralized MGST algorithm does show an advantage when the graph state is relatively dense (with a high average degree), as shown in Fig. 7 (c), which is understandable due to the fact that it treats vertices instead of edges as ’files’ and shall be more efficient when the average number of edges is high and grows quadratically with the number of vertices. However, we remark here that there can be the case that P2PGSD series algorithms offer better solutions even when the graph state is a complete graph, and we offer an example in the Appendix. A.1.
@@
Varying channel probabilities. The robustness under the probabilistic network of algorithms is investigated under various average channel probabilities assuming distribution of the star state and the grid state, the results of which are shown in Fig. 8. As we expected, P2PGSD-Max has the best performance in terms of shots when the probabilities decrease, since we are trading memory for speed. In fact, P2PGSD-Max also tends to consume less memory than other algorithms when the probability decreases. It is a well-known fact that saving all the unused bell pairs in a path can result in a better scaling of the expected time for success (Azuma et al., 2023), and this will lead to the threshold behavior in resource consumption, which helps explain the result being observed. Since this is a path-level optimization and can be considered as part of path recovery, we did not implement it in our main distribution algorithms; instead, we offer an analytic solution to the threshold probability for some simple cases in the Appendix B. In particular, for the grid state distribution case shown in Fig. 8 (b), MGST seems to perform better than P2PGSD when the probability is below 0.5. In addition, both ST version algorithms perform increasingly better than P2PGSD when probability decreases. It is also worth noting that although the factor version of ST-P2PGSD does not perform well when the probability is high, which may be due to the finite precision effect in calculating the metric, it gains significant advantages in both shot consumption and cumulative memory consumption when the probability decreases. A similar threshold behavior is also observed when comparing its performance with ST-P2PGSD shown in Fig. 8.
@@
answer earlier. All P2P-based algorithms show similar scaling with an exponential factor of 2.5 while MGST shows a cubic scaling when increasing the number of nodes in the network, this is somehow expected for the P2PGSD series algorithms due to the nonlinear increasing of the edges. However, MGST performs far better than its upper bound. This may be due to the fact that the maxflow algorithm may have an average run-time scaling close to $\omega(|E_{N}|)$ instead of its worse upper bound $\mathscr{O}(|V_{N}|^{2}|E_{N}|)$ due to the constant maximum flow in our graph state-distributing task.

### v3

@@ Abstract.
Graph states are a class of important multiparty entangled states, of which Bell pairs are the special case. Realizing a robust and fast distribution of arbitrary graph states in the downstream layer of the quantum network can be essential for further large-scale quantum networks. We propose a novel quantum network protocol called P2PGSD inspired by the classical Peer-to-Peer (P2P) network to efficiently implement the general graph state distribution in the network layer, which demonstrates advantages in resource efficiency and scalability over existing methods for sparse graph states. An explicit mathematical model for a general graph state distribution problem has also been constructed, above which the intractability for a wide class of resource minimization problems is proved and the optimality of the existing algorithms is discussed. In addition, we leverage the spacetime quantum network inspired by the symmetry from relativity for memory management in network problems and use it to improve our proposed algorithm. The advantages of our protocols are confirmed by numerical simulations showing an improvement of up to 50% for general sparse graph states, paving the way for a resource-efficient multiparty entanglement distribution across any network topology.
@@ 2.2. Principles of P2P Graph State Distribution
Here, $G_{N}=(V_{N},E_{N})$ is an undirected graph representing the topology of the network with $V_{N}$ denoting the set of nodes and $E_{N}$ denoting the set of channels. $W_{c}:E_{N}\rightarrow\mathbb{N}^{+}$ is a map that represents the width of the channel such that each channel $c\in E_{N}$ can establish $W_{c}(c)$ Bell pairs per cycle time. And $W_{m}:V_{N}\rightarrow\mathbb{N}$ is a map that represents the long-term memory available for each node.
@@ 2.3. Problem Formulation
An One-shot Bell Pair Strategy $b_{1}$ is a map $b_{1}:\vec{E}_{N}\times V_{S}\rightarrow\{+1,0,-1\}$ with 00 representing that the channel is not used by the corresponding vertex and it fulfills the following flow conditions:
@@
    $\exists n_{j}(\neq n_{i})\in N_{v_{s}}^{\mathscr{b}_{N}},\,b_{k}\in\mathscr{b}_{N},\vec{c}_{ij}=(c_{ij},n_{j},n_{i})\in\vec{E}_{N},\mathrm{s.t.}\,b_{k}(\vec{c}_{ij},v_{s})>0$.
@@
A $N$-Shot Bell Pair Flow Strategy $[\mathscr{b}_{N}]$ is an equivalent class defined by the following equivalent relation for $N$-shot Trial Solution as
$\mathscr{b}_{N}\sim\mathscr{b}^{\prime}_{N}\Leftrightarrow\sum_{b_{i}\in\mathscr{b}_{N}}{|b_{i}(\vec{e},v_{s})|}=\sum_{b^{\prime}_{i}\in\mathscr{b}^{\prime}_{N}}{|b^{\prime}_{i}(\vec{e},v_{s})|},\,\forall v_{s}\in V_{S},\,\forall\vec{e}\in\vec{E}_{N}.$
@@ 2.4. Optimizing Objectives
The first metric is simply the number of cycles required, and it is also the time consumption to execute such a solution. The second is frequently considered in the previous literature (Cuquet and Calsamiglia, 2012; Pirker et al., 2018; Meignant et al., 2019; Fischer and Towsley, 2021; Xie et al., 2021; Fan et al., 2024), while the Long-term Memory Consumption, as we explained in the previous subsection, is just the cumulative memory (Alwen and Serbinenko, 2015; Beame and Kornerup, 2023) and can be considered as a counterpart of the second. We remark that some qubits during execution can be measured instantaneously, while some other qubits will be required to be kept to the next shot. The latter must be kept for at least one cycle time, where the consumption of quantum memory is necessary and the usage is counted by our defined metric.
@@ 2.5. Complexity Analysis
The aforementioned theorem is important in the sense that it forces and also rationalizes it to pursue a heuristic algorithm that consumes low resources with relatively low runtime, since it is nearly impossible to have a scalable and practical one that converges to the global minimum unless $P=NP$. However, we get a tight upper bound for #$\#\mathscr{S}$ in the case where there is unlimited memory. Here, ”tight” means that we construct an instance $\mathscr{P}$ that requires at least the upper bound number of shots to distribute using any algorithm. This result will help us analyze the performance of the proposed algorithms in the worst case and also ensures that the algorithms obtained are within a polynomial approximation (linear to optimize #$\#\mathscr{S}$) and thus reliable.
@@ 3.1. Network Model and Routing Protocol
    The high-quality Bell pair is created in a heralded manner (it can be an abstraction of the physical devices that generate high-quality Bell pairs from multiple tries and entanglement purification, including the quantum repeater(Bennett et al., 1996; Bernien et al., 2013; Pirandola et al., 2017)). A channel of width $W_{c}(e)$ with $e\in E_{N}$ is capable of creating at most $W_{c}(e)$ Bell pairs within a time interval, each with a probability of $P_{c}(e)$.
@@
    All single-qubit operations are assumed to be deterministic, but the only two-qubit operation, $CZ$ operation, is considered probabilistic with probability $P_{s}$ but in a heralded way (Li et al., 2020; Sung et al., 2021).
@@ 3.2. Modified Graph State Transfer Algorithm
Figure 3. Illustration for MGST algorithm. (a) The constructed flow graph with the subgraphs inside the red shadow box represents one of the copies of $G_{N}$ labeled as $G_{N}^{i},\,i=1...k$ together with virtual nodes (larger circles) and virtual arcs (dotted arrows) with the numbers in different colors representing the width of the corresponding channel. (b) An example of $3$-shot solution of MGST. The nodes with multiple colors are the root. The paths in different colors represent the edge-disjoint spacetime path from $\{(\alpha_{D}(v_{s}),4\}$ to the root node $(n_{\mathrm{root}},1)$. The blue dotted arcs show the possible recovery possible if the current path fails.
@@
The GST algorithm proposed by Fischer et al. (Fischer and Towsley, 2021) is an efficient algorithm but it only gives us an equivalent class $[\mathscr{b}_{N}]$ that solves $\mathscr{P}$. We need to pick an element $\mathscr{b}_{N}$ from $[\mathscr{b}_{N}]$ as well as find a $\mathscr{m}_{n}$ to form the solution $\mathscr{S}_{N}=\{\mathscr{b}_{N},\mathscr{m}_{N}\}$ so that our network can execute the solution. One can certainly use the flow decomposition theorem to decompose $[\mathscr{b}_{N}]$ into a set of paths; however, it is unknown whether such paths can be executed in $N$ shots without conflicting or taking apart specific paths into several shots (which we call it simply executable) is unknown. In the worst case, we may have to consume a lot of memory to execute a single path part-by-part across multiple shots, which is definitely unwanted. Hence, how to decompose the flow into a $N$-shot simple executable path set is not a trivial problem. We propose the following novel construction of the flow graph to find the $|V_{N}|$ edge-disjoint paths from a given root $n_{\mathrm{root}}\in V_{N}$ executable in $k$ shots:
@@
However, there is one question left to be solved: Will finding such a simple executable path set results in the use of more shots than the flow solution given by GST. This question is answered by the following theorem proven in Appendix. A.1.
@@ 3.3. Peer-to-Peer Graph State Distribution Algorithm
Figure 4. Illustration for P2PGSD algorithm. (a) When the path between $n_{v_{s_{1}}}$ (yellow) and $n_{v_{s_{2}}}$ (blue) is first found, the assignment of the Bell pairs (yellow & blue) is not determined. After the new path to $n_{v_{s_{3}}}$ (green node) is found from an ’unfixed’ node along Path 1, part of the Path is fixed and it is split into the new Path 1 and Path 2, while Path 3 is not fixed (yellow & green). (b) The P2PGSD solution under Minimum memory strategy. The half-and-half colors indicate the node that implements the edge for the corresponding vertices. (c) The P2PGSD solution under Standard memory strategy. (d) The P2PGSD solution under Maximum memory strategy. This might be helpful for later trials (dotted arcs) in case the previous trial fails.
@@
We omitted the proof of the correctness for this simple modification version of Dijkstra, while we remark that the correctness of this algorithm relies on the assumption that in the current network, no path from other $n_{v_{s_{1}},i}/n_{v_{s_{2}},j}$ to other $n_{v_{s_{1}},i^{\prime}}/n_{v_{s_{2}},j^{\prime}}$ with smaller cost exits is available. The modified version of the Dijkstra algorithm has a time complexity of $O(|E_{N}||V_{N}|)$ where the additional $|V_{N}|$ factor comes from the fact that we need to calculate the metric for the whole path at each time to allow a more general metric beyond the addable one, though we can turn our metric into the latter in this work.
@@
    Maximum: All the nodes’ ancillary vertices with ’fixed’ flag will be kept to the next shot.
@@ 3.4. Spacetime Peer-to-Peer Graph State Distribution Algorithm
Figure 5. Illustration for STP2PGSD algorithm. (a) A no-plan-ahead solution compared to (b) a solution without such limitation, where the parallel arcs in the shot $k$ of the lower plane indicate the violation of such limitation. (b) STP2PGSD directly finds a path in the spacetime network and overcomes the limitation of no-plan-ahead. The addition vertical arcs represent a recovery mechanism similar to MGST.
@@
The pseudocode for the STP2PGSD algorithm is shown in Alg. 3 and the runtime of STP2PGSD is upper bound by $O(|V_{S}|^{3}|E_{S}||V_{N}||E_{N}|\log|V_{S}|)$.
@@ 3.5. Adaptive Protocol
At each shot, only the plan for the first shot given by the algorithm will be collected and executed by the network. After that, the execution results for the current shot will be collected in the synchronization phase and the distribution task $\mathscr{P}$ will be modified accordingly. Then the modified distribution task together with the available resources from the last shot will be submitted to the algorithm in the next cycle to perform an adaptive distribution strategy. This mechanism offers strong flexibility for the graph state distribution in a probabilistic network and also inherits a sense of recovery mechanism using the link resources from the previous shot, which will be beneficial to scaling issues when the probabilities of success for the links in the network are low. However, such an adaptive protocol will require some modifications to the original algorithm to take care of the resources from the previous shot, which is easy and trivial for the first two algorithms but requires careful design for the ST-P2PGSD algorithm. Further details of the adaptive protocol will be left for discussion in the Appendix. C.
@@ 4.1. Simulation Setup
The proposed network models as well as the algorithms are implemented on a custom-built simulator in Python with topology generation and common subroutines adapted from the NetworkX package (Hagberg et al., 2008). As we have mentioned in Sec. 3.1, the protocols presented can be considered a network protocol that regards general graph state as a basic connection instead of Bell pairs. Thus, our simulations are not built upon other proposed protocols, but rather be implemented from the relatively basic level. The source codes as well as the data for the numerical evaluation can be found in the corresponding open repository (Huang_P2PGraphStatesDistribution).
@@
Multiple types of graph states are being considered including the Erdős-Rényi graph (ERDdS and R&wi, 1959), star graphs, Prüfer tree graphs (Wang et al., 2009), grid graphs, etc., with some of them being generated randomly when possible. After the graph state topology is generated, the vertices are randomly assigned to the network nodes to implement the assignment map $\alpha_{D}$.
@@ 4.2. Metrology
The simulation is designed to evaluate the time shots, #$\#\mathscr{S}$, and cumulative memory consumption, $\#\mathscr{M}$, defined in Sec. 2.2 as well as the computation runtime for distributing various types of graph states (random or fixed) on a random network under the algorithms proposed in this work as well as the one modified from previous work (MGST). We do not focus on Bell pair usage, $\#\mathscr{B}$, since it is not our main optimization goal and it can be shown that it is not compatible with the optimization of the previous two resources and requires a completely different design strategy for the algorithms.
@@ 4.3. Results Analysis
The general performance of protocols based on 3 main algorithms together with two additional variants for P2PGSD and ST-P2PGSD are shown in Fig. 7, where we conclude that the developed P2P series algorithms demonstrate a huge advantage over the center-based MGST algorithm in distributing the widely used tree graphs and grid graphs in the sense of shot usages and cumulative memory usage. The resource consumption difference grows when the number of vertices of the graph state increases, which supports our statement that many resources can be wasted if the graph state topology is not considered for distribution. Furthermore, P2PGSD with a maximum memory strategy consumes more memory than P2PGSD while producing almost the lowest shot consumption, which illustrates the idea of trading memory for faster distribution. It is observed that the memory usage of ST-P2PGSD falls between that of P2PGSD and P2PGSD-Max, and similarly, its shot consumption is intermediate between the two as well. We also investigate the performance of the algorithms under extremely limited memory as shown in Fig. 7(c) where each node only holds the amount of long-term memory that is equal to an average memory of 1 or its assigned vertices to fulfill the minimum requirement to distribute such a graph state (while one of the nodes in the network is chosen to possess $2|V_{S}|$ memory in order for MGST to be executed). In this case, MGST consumes significantly more resources than the P2PGSD algorithms series and also has huge variance due to the available center node choice. In addition, the performance of various P2PGSD algorithms is drawn to a common line since there is not much room for memory optimization.
@@
Figure 7. Shots (bar) and cumulative memory (scatter) usage of protocols based on various algorithms versus the number of vertices in certain types of graph state. P2PGSD adapts the standard memory strategy while P2PGSD-Max adapts the maximum memory strategy. The shaped area and the error bar represent the standard deviation of the data. The metric of standard ST-P2PGSD is given by Eq. (12) while the ”Factor” version is given by Eq. (11). (a) Random Prüfer tree graphs. (b) Grid graphs. (c) Grid graphs while the performance is evaluated under extremely limited memory. (d) Star graphs. (e) Bell pairs (regular 1) graphs with cell topology network (illustrated in subfigure with width 1 and average channel probability $0.9$). (f) Erdős-Rényi with average edge probability $0.1$.
@@
Advantageous tasks. Three different main algorithms establish significant advantages with respect to distribution tasks with certain topology characteristics. Fig. 7 (d)-(f) shows the typical performance comparison between the algorithms on the presentative tasks. As our initial inspiration for this work, the (ST-)P2PGSD series algorithms demonstrate extremely low resource consumption when the state being distributed is a star graph (GHZ state), as shown in Fig. 7 (d). One observes that shot usage as well as cumulative memory consumption hardly grows under the P2PGSD series algorithms, in contrast to MGST. This is a straightforward consequence since distributing such a state only requires the assigned node to be connected in tree form by the successful Bell pair and the resource consumption saturates if the network size is unchanged due to the deterministic local operation. The spacetime version algorithms do not show any advantages in the previous case since these tasks admit a one-shot solution and do not need memory optimization, but one observes from Fig. 7 (b) that they consume remarkably fewer resources in terms of both shot usages and memory usages when the network suffers from serious discontinuous bottleneck channels as the case of the so-called ”cell topology network” shown in Fig. 7 (g). The ST-network method makes it rather easy and convenient to deal with this problem and generate an efficient memory management plan. Finally, the centralized MGST algorithm does show an advantage when the graph state is relatively dense (with a high average degree), as shown in Fig. 7 (c), which is understandable due to the fact that it treats vertices instead of edges as ’files’ and shall be more efficient when the average number of edges is high and grows quadratically with the number of vertices. However, we remark here that there can be the case that P2PGSD series algorithms offer better solutions even when the graph state is a complete graph, and we offer an example in the Appendix. A.1.
@@
Varying channel probabilities. The robustness under the probabilistic network of algorithms is investigated under various average channel probabilities assuming the distribution of the star state and the grid state, the results of which are shown in Fig. 8. As we expected, P2PGSD-Max has the best performance in terms of shots when the probabilities decrease, since we are trading memory for speed. In fact, P2PGSD-Max also tends to consume less memory than other algorithms when the probability decreases. It is a well-known fact that saving all the unused Bell pairs in a path can result in a better scaling of the expected time for success (Azuma et al., 2023), and this will lead to the threshold behavior in resource consumption, which helps explain the result being observed. Since this is a path-level optimization and can be considered as part of path recovery, we did not implement it in our main distribution algorithms; instead, we offer an analytic solution to the threshold probability for some simple cases in the Appendix B. In particular, for the grid state distribution case shown in Fig. 8 (b), both ST version algorithms perform increasingly better than P2PGSD when probability decreases. It is also worth noting that although the factor version of ST-P2PGSD does not perform well when the probability is high, which may be due to the finite precision effect in calculating the metric, it gains significant advantages in both shot consumption and cumulative memory consumption when the probability decreases. A similar threshold behavior is also observed when comparing its performance with ST-P2PGSD shown in Fig. 8.
@@
answer earlier. All P2P-based algorithms show similar scaling with an exponential factor of 2.5 while MGST shows a cubic scaling when increasing the number of nodes in the network, this is somehow expected for the P2PGSD series algorithms due to the nonlinear increasing of the edges. However, MGST performs far better than its upper bound. This may be due to the fact that the maxflow algorithm may have an average run-time scaling close to $\omega(|E_{N}|)$ instead of its worse upper bound $O(|V_{N}|^{2}|E_{N}|)$ due to the constant maximum flow in our graph state-distributing task.